# The NeRFect Match: Exploring NeRF Features for Visual Localization

### v1

@@ Abstract
In this work, we propose the use of Neural Radiance Fields (NeRF) as a scene representation for visual localization. Recently, NeRF has been employed to enhance pose regression and scene coordinate regression models by augmenting the training database, providing auxiliary supervision through rendered images, or serving as an iterative refinement module. We extend its recognized advantages – its ability to provide a compact scene representation with realistic appearances and accurate geometry – by exploring the potential of NeRF’s internal features in establishing precise 2D-3D matches for localization. To this end, we conduct a comprehensive examination of NeRF’s implicit knowledge, acquired through view synthesis, for matching under various conditions. This includes exploring different matching network architectures, extracting encoder features at multiple layers, and varying training configurations. Significantly, we introduce NeRFMatch, an advanced 2D-3D matching function that capitalizes on the internal knowledge of NeRF learned via view synthesis. Our evaluation of NeRFMatch on standard localization benchmarks, within a structure-based pipeline, sets a new state-of-the-art for localization performance on Cambridge Landmarks.
@@ 1 Introduction
NeRF has been leveraged for tackling visual localization in various ways. iNeRF-style approaches [73, 18] utilize a pre-trained NeRF as an inference-time pose refinement module. Yet such methods commonly suffer from slow convergence and require pose initialization to be provided. The end-to-end APR and SCR methods use a pre-trained NeRF only at train-time to augment training samples [43, 15], provide consistency supervision [16, 17], and generate proxy depth ground-truth [15]. In this case, NeRF merely serves as an auxiliary representation. Orthogonal to the above ways of leveraging NeRF, recent works [38, 44] propose to use NeRF as a flexible 3D model that can be enriched with volumetric descriptors to establish 2D-3D matches between an image and the scene. CrossFire [44] augments the base NeRF model with an additional 3D feature prediction branch. This feature is supervised to match 2D features extracted by an image backbone. NeRF-Loc[38] performs feature matching by utilizing a combination of features extracted from a generalizable NeRF and projected multi-view image features as 3D point features. However, both methods require training NeRF jointly with the matching task, prohibiting the usage of pre-built NeRF scenes.
In contrast, our work treats NeRF as the primary scene representation in visual localization without re-training and modifications. Specifically, we focus on a crucial component – NeRF features – and (i) demonstrate their inherent capability in effectively supporting feature matching. (ii) We introduce, NeRFMatch, a matching transformer that aligns 2D image features with 3D NeRF features and a minimal version of it to facilitate real-time applications. (iii) We present two options for performing pose refinement on top of NeRFMatch and conduct detailed analysis on their refinement effectiveness. (iv) We use our NeRFMatch and pose refinement modules to perform hierarchical NeRF localization which achieves the state-of-the-art localization performance on Cambridge Landmarks [32]. Based on our experiments, we point out future work in needs to improve our indoor localization performance. Our research paves the path towards localization leveraging NeRF as the sole representation of the scene.
@@ 3.2 NeRF for Localization
NeRF architecture.  NeRF [42] models a 3D scene with a coordinate network. Specifically, it maps a 3D point and a camera viewing direction to their corresponding volume density and RGB values. As depicted in Fig. 2, a standard NeRF model consists of two non-learnable positional encodings, $P_{x}$ for 3D coordinates and $P_{d}$ for viewing directions, and three trainable components: a 3D point encoder $\Theta_{x}$, a volume density decoder $\Theta_{\sigma}$, and a RGB color decoder $\Theta_{c}$. More concretely, the 3D encoder consists of $L$ intermediate layers, i.e., $\Theta_{x}=\Theta^{L}_{x}\circ\dots\circ\Theta^{1}_{x}$. In this work, we are particularly interested in exploring the potential of 3D features extracted within the 3D encoder. Given a 3D point $X\in R^{3}$, we define its 3D feature extracted at $j$-th 3D encoder layer as $f^{j}=\Theta^{j}_{x}\circ\dots\circ\Theta^{1}_{x}(P_{x}(X))$, where $\Theta^{j}_{x}$ is the $j$-th layer in the 3D encoder. The last layer 3D feature is next input into the density decoder to predict the volume density $\sigma=\Theta_{\sigma}(f^{L})$, while the color decoder takes a positional-embedded viewing direction $d\in R^{2}$ in addition to the 3D feature $f^{L}$ to compute the view-dependent color $c=\Theta_{c}(f^{L},P_{d}(d))$.
@@
Here, $X_{i}$ is $i$-th sampled 3D points, $f^{j}_{i}$ is its 3D feature extracted at $j$-th 3D encoder layer and $w_{i}$ is the corresponding weight computed from its volume density prediction as defined in Eq. 1.
@@ 3.3 Pose Refinement
Following the matching-based pose estimation, we employ two approaches to refine the estimated camera pose. The first approach, iterative refinement, uses the estimated camera pose as a new reference for extracting NeRF features. This process involves repeating the matching procedure with the updated reference pose, incrementally enhancing the results due to the closer proximity of the reference pose, which makes the NeRF and image features more similar. While this process can be repeated multiple times, significant improvements are typically observed after the first refinement (cf. Sec. 5.4). Inspired by iNeRF [73], our second refinement option combines optimization and matching where we backpropagate through the frozen NeRF model to optimize an initial camera pose estimate by minimizing the photometric difference between the query image and the current NeRF rendered image. We then use the optimized camera pose to render again 3D points and features to perform the 2D-3D matching and compute the final refined camera pose. In this option, the iNeRF-style optimization is utilized as an intermediate refinement module allowing to find 3D points and features that lead to improved matching.
@@ 4 Image-to-NeRF Matching Network
Figure 3: NeRFMatch architecture. We present NeRFMatch as our full matching model (rightmost) and NeRFMatch-Mini as a light version of it (middle). Both models share the same feature extraction process, where we use a 2D encoder to extract image features at two resolutions and render 3D points with associated NeRF features at sampled 2D pixel locations from the reference viewpoint.
Overview.  To establish 2D-3D correspondences between a query image and NeRF scene points for localization, we introduce two variants of our proposed image-to-NeRF matching network, a full version (NeRFMatch) and a minimal version (NeRFMatch-Mini), as depicted in Fig. 3. The full version is notably more expressive, incorporating powerful attention modules [52, 59, 14] and follows a coarse-to-fine matching paradigm [59, 77, 14]. While it delivers more accurate matches it is computationally more expensive (cf. Sec. 5.3). Consequently, we also propose a minimal version of our method that focuses on learning good features for matching, removing the need to learning the matching function itself. Both models comprise a feature extraction module that encodes both the query image and 3D scene points into a feature space, and a matching module that aligns these two feature sets to determine the 2D-3D correspondences. We further detail their architectural designs and supervision mechanisms in the subsequent subsections.
@@ 5 Experiments
Implementation details.  We use the first two blocks of ConvFormer [74] as the image backbone and initialize it with ImageNet-1K [50] pre-trained weights [^1] The weights can be downloaded from huggingface.co/timm/convformer_b36.sail_in1k_384. We set feature dimensions for coarse and fine matching as $D^{c}=256$ and $D^{f}=128$. For fine matching, we use local window size $w=5$ for image feature cropping. We resize query images to $480\times 480$ for all experiments. We train minimal and full NeRFMatch models using Adam [33] optimizer with canonical learning rate $clr=0.0008/0.0004$ and batch size $cbs=16$ for $30/50$ epochs accordingly. We decay the learning rate based on the cosine annealing scheduling [39]. Our models are trained on $8$ Nvidia V100 GPUs (16/32GB). A MipNeRF model is $5.28MB$ in size. Given a camera pose, it takes $141$ milliseconds to render $6003$D points with its features on a single 16GB Nividia V100 GPU. We provide NeRF implementation details in the supplementary material.
@@
NeRFMatch-Mini NeRF+RGB 11.2/0.3 19.3/0.6 5.2/0.5 11.2/0.4 30.6/0.2 15.5/0.4
NeRFMatch NeRF+RGB 6.4/0.2 10.3/0.4 6.2/0.5 7.8/0.3 22.0/0.1 10.5/0.3
NeRFMatch NeRF 6.4/0.2 9.4/0.4 6.2/0.4 11.9/0.4 23.1/0.1 11.4/0.3
5.1 Localization Evaluation
Inference settings.  For both NeRFMatch models, we retrieve its top-1 reference pose for 3D rendering and render NeRF features from the 3rd encoder layer and their best pose refinement setting validated by our ablation studies.
Results on Cambridge Landmarks.  As shown in Tab. 1, our minimal version despite being lightweight, is able to achieve competitive results w.r.t. most of the SCR methods and surpass all APR methods. With a more advanced attention-based hierarchical matching function, our full model achieves the state-of-the-art results among all methods. Our experiments fully demonstrate that the NeRF inner features learned via view synthesis are discriminative 3D representations for 2D-3D matching.
@@ 5.1 Localization Evaluation
NeRFMatch-Mini NeRF+RGB 1.4/0.5 1.7/1.0 2.1/0.7 4.4/1.0 4.7/1.0 2.2/0.5 8.8/2.1 3.6/0.9 67.9
NeRFMatch NeRF+RGB 0.9/0.3 1.3/0.4 1.6/1.0 3.2/0.7 3.3/0.7 1.3/0.3 7.5/1.3 2.7/0.7 75.3
NeRFMatch NeRF 0.9/0.3 1.3/0.4 1.6/1.0 3.3/0.7 3.2/0.6 1.3/0.3 7.2/1.3 2.7/0.7 75.4
Table 2: Indoor localization on 7-Scenes [57]. We report per-scene median rotation and position errors in $(cm,^{\circ})$ and its average across scenes. We further report the averaged localization recall.
@@
NeRF-only localization.  We further push our method to using NeRF as the only scene representation for localization, which means we would no longer need access to a real image database as part of the scene representation, which requires a significantly bigger storage than a single NeRF model. For this purpose, we propose to perform image retrieval on synthesized images rendered by our NeRF model. This entails a small decrease in performance on Cambridge, e.g., $0.9cm$ in translation error. This slight degradation in performance can be attributed to the increased complexity of these scenes for NeRF. This assertion is supported by the observation that, when switching to image retrieval on synthesized images, our model demonstrates virtually no change in performance on the indoor 7-Scenes dataset. Our results open the door to a localization pipeline where only a NeRF model is needed.
@@ 5.2 NeRF Feature Ablation
Med. Translation ($cm,\downarrow$) 432.3 25.5 22.9 23.3 21.8 22.3 23.5 24.1 40.9
Med. Rotation (${}^{\circ},\downarrow$) 6.5 0.6 0.5 0.5 0.5 0.5 0.5 0.5 1.0
@@
Table 3: NeRF feature ablation on Cambridge Landmarks [32]. We train NeRFMatch-Mini with different 3D features and compare their localization performance.
@@
- ResNet34 Minimal Per-Scene 24.8/0.6 61.5 32.8 23
- ConvFormer Minimal Per-Scene 26.3/0.6 60.8 42.8 37
NeRFMatch-Mini PS ConvFormer Minimal Per-Scene 21.8/0.5 65.8 42.8 37
NeRFMatch-Mini MS ConvFormer Minimal Multi-Scene 23.9/0.5 62.1 42.8 37
NeRFMatch PS ConvFormer Full Per-Scene 13.1/0.3 81.4 50.4 157
NeRFMatch MS ConvFormer Full Multi-Scene 16.0/0.4 76.8 50.4 157
5.3 NeRFMatch Ablation
Training ablation.  In addition to architecture choices, we also examine the influence of different training settings, i.e., training per-scene and training multi-scenes (all scenes within a dataset). Despite NeRF features being trained per-scene, we surprisingly find that both minimal and full NeRFMatch models can be trained to handle multi-scene localization with only a slight decrease in accuracy w.r.t. the per-scene model. This is a similar finding as recent scene-agnostic SCR methods [72, 63, 62], that extend per-scene SCR to multi-scenes by conditioning SCR on scene-specific 3D points. While scene-agnostic SCR learns to regress directly the 3D features in the form of xyz coordinates, our models learn to find a common ground between image features and NeRF features for matching.
@@ 5.4 Pose Refinement
Model Init Iter-1 Opt-1+Match
NeRFMatch-Mini PS 21.8/0.5/65.8 21.0/0.5/67.9 15.5/0.4/77.9
NeRFMatch-Mini MS 23.9/0.5/62.1 23.2/0.5/64.6 16.6/0.4/75.6
NeRFMatch PS 13.1/0.3/81.4 10.5/0.3/87.1 11.6/0.3/85.1
NeRFMatch MS 16.0/0.4/76.8 13.7/0.3/82.3 14.4/0.4/80.1
Figure 4: Refinement ablation on Cambridge Landmarks [32]. On the left side, we depict the average recall for optimization-based (Opt+Match) and iterative (Iter) refinement approaches across multiple iterations. We provide results for both the NeRFMatch and its minimal setting, with "MS" denoting multi-scene trained and "PS" indicating per-scene trained models. On the right side, we report averaged median pose error in ($cm/^{\circ}$) and localization recall with a single refinement iteration.
@@
After conducting matching architecture ablation, we investigate different refinement methods to further increase pose accuracy. We examine two approaches: (i) an iterative approach, where we re-run matching with the last computed query estimation as reference pose, and (ii) an optimization approach, where we optimize a reference pose through frozen NeRF weights using a photometric loss and then run the final matching. As shown in Fig. 4, both methods show an improvement over the initial estimate. We assess both the NeRFMatch model and its minimal version (Mini). For each version, we evaluate models trained per-scene (PS) and models trained on multiple scenes (MS). The most substantial improvement for both approaches occurs in the first iteration, after which the estimates tend to stabilize.
@@
Iterative refinement for computational efficiency.  In the NeRFMatch setting, both refinement approaches show similar results, since the initial pose estimate is relatively close to the solution. However, it is worth noting that the optimization approach incurs a higher computational cost compared to the iterative approach. The runtime for the optimization refinement excluding the matching step (shown in Tab. 4, depending on the model) is $398.4ms$ for a single optimization step and subsequent rendering. In contrast, for a single step of iterative refinement, the runtime amounts to $141.2ms$. Given the slightly better performance and the quicker computational time, we opt for the iterative approach as the default refinement for NeRFMatch.
@@
Optimization refinement for large pose corrections.  In comparison to the iterative method, the optimization approach significantly benefits the minimal matching model. This is because the initial query pose estimation is notably distant compared to the one given by NeRFMatch, and subsequent iterations only result in incremental refinements. The optimization approach takes a more substantial step towards the query pose, and achieving a more optimal reference pose results in improved query pose estimation. We have already seen that the optimization refinement incurs higher computational time, but it also depends on a learning rate schedule, which is sensitive to configuration. Through empirical analysis, we select $1\times 10^{-5}$ as the initial learning rate for NeRFMatch and $0.001$ for minimal settings, with exponential decay. We set the optimization approach as the default refinement for NeRFMatch-Mini.

### v2

# https://nerfmatch.github.io

@@ Abstract
In this work, we propose the use of Neural Radiance Fields (NeRF) as a scene representation for visual localization. Recently, NeRF has been employed to enhance pose regression and scene coordinate regression models by augmenting the training database, providing auxiliary supervision through rendered images, or serving as an iterative refinement module. We extend its recognized advantages – its ability to provide a compact scene representation with realistic appearances and accurate geometry – by exploring the potential of NeRF’s internal features in establishing precise 2D-3D matches for localization. To this end, we conduct a comprehensive examination of NeRF’s implicit knowledge, acquired through view synthesis, for matching under various conditions. This includes exploring different matching network architectures, extracting encoder features at multiple layers, and varying training configurations. Significantly, we introduce NeRFMatch, an advanced 2D-3D matching function that capitalizes on the internal knowledge of NeRF learned via view synthesis. Our evaluation of NeRFMatch on standard localization benchmarks, within a structure-based pipeline, achieves competitive results for localization performance on Cambridge Landmarks. We will release all models and code.
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
NeRF has been leveraged for tackling visual localization in various ways. iNeRF-style approaches [73, 16] utilize a pre-trained NeRF as an inference-time pose refinement. Yet such methods commonly suffer from slow convergence and require pose initialization to be provided. The end-to-end APR and SCR methods use a pre-trained NeRF only at train-time to augment training samples [44, 15], provide consistency supervision [18, 17], and generate proxy depth ground-truth [15]. In this case, NeRF merely serves as an auxiliary representation. Orthogonal to the above ways of leveraging NeRF, recent works [38, 43] propose to use NeRF as a flexible 3D model that can be enriched with volumetric descriptors to establish 2D-3D matches between an image and the scene. CrossFire [43] augments the base NeRF model with an additional 3D feature prediction branch. This feature is supervised to match 2D features extracted by an image backbone. NeRF-Loc[38] performs feature matching by utilizing a combination of features extracted from a generalizable NeRF and projected multi-view image features as 3D point features. However, both methods require training NeRF jointly with the matching task, prohibiting the usage of pre-built NeRF scenes.
In contrast, our work treats NeRF as the primary scene representation in visual localization without re-training and modifications. Specifically, we focus on a crucial component – NeRF features – and (i) demonstrate their inherent capability in effectively supporting feature matching. (ii) We introduce, NeRFMatch, a matching transformer that aligns 2D image features with 3D NeRF features and a minimal version of it to facilitate real-time applications. (iii) We present two options for performing pose refinement on top of NeRFMatch and conduct detailed analysis on their refinement effectiveness. (iv) We use our NeRFMatch and pose refinement modules to perform hierarchical NeRF localization which achieves competitive localization performance on Cambridge Landmarks [32]. Based on our experiments, we point out future work in needs to improve our indoor localization performance. Our research paves the path towards localization leveraging NeRF as the sole representation of the scene.
@@ 3.2 NeRF for Localization
NeRF architecture.  NeRF [42] models a 3D scene with a coordinate network. Specifically, it maps a 3D point and a camera viewing direction to their corresponding volume density and RGB values. As depicted in Fig. 2, a standard NeRF model consists of two non-learnable positional encodings, $P_{x}$ for 3D coordinates and $P_{d}$ for viewing directions, and three trainable components: a 3D point encoder $\Theta_{x}$, a volume density decoder $\Theta_{\sigma}$, and a RGB decoder $\Theta_{c}$. More concretely, the 3D encoder consists of $L$ layers, i.e., $\Theta_{x}=\Theta^{L}_{x}\circ\dots\circ\Theta^{1}_{x}$. In this work, we are particularly interested in exploring the potential of 3D features extracted within the 3D encoder. Given a 3D point $X\in R^{3}$, we define its 3D feature extracted at $j$-th 3D encoder layer as $f^{j}=\Theta^{j}_{x}\circ\dots\circ\Theta^{1}_{x}(P_{x}(X))$, where $\Theta^{j}_{x}$ is the $j$-th layer in the 3D encoder. The last layer 3D feature is next input into the density decoder to predict the volume density $\sigma=\Theta_{\sigma}(f^{L})$, while the color decoder takes a positional-embedded viewing direction $d\in R^{2}$ in addition to the 3D feature $f^{L}$ to compute the view-dependent color $c=\Theta_{c}(f^{L},P_{d}(d))$.
@@
Here, $X_{i}$ is $i$-th sampled 3D points, $f^{j}_{i}$ is its 3D feature extracted at $j$-th 3D encoder layer and $w_{i}$ is the weight computed from its density prediction Eq. 1.
@@ 3.3 Pose Refinement
Following the matching-based pose estimation, we employ two approaches to refine the estimated camera pose. The first approach, iterative refinement, uses the estimated camera pose as a new reference for extracting NeRF features. This process involves repeating the matching procedure with the updated reference pose, incrementally enhancing the results due to the closer proximity of the reference pose, which makes the NeRF and image features more similar. While this process can be repeated multiple times, significant improvements are typically observed after the first refinement (cf. Sec. 5.4). Inspired by iNeRF [73], our second refinement option combines optimization and matching where we backpropagate through the frozen NeRF model to optimize an initial camera pose estimate by minimizing the photometric difference between the query image and the current NeRF rendered image. We then use the optimized camera pose to render again 3D points and features to perform the 2D-3D matching and compute the final refined camera pose.
@@ 4 Image-to-NeRF Matching Network
Figure 3: NeRFMatch architecture. We present NeRFMatch as our full matching model (rightmost) and NeRFMatch-Mini as a light version of it (middle). Both models share the same feature extraction process, where we use a 2D encoder to extract image features at two resolutions and render 3D points with associated NeRF features at sampled 2D pixel locations from the reference viewpoint. The full matching uses self-attention (SA) and cross-attention (CA) with positional encodings (PE).
Overview.  To establish 2D-3D correspondences between a query image and NeRF scene points for localization, we introduce two variants of our proposed image-to-NeRF matching network, a full version (NeRFMatch) and a minimal version (NeRFMatch-Mini), as depicted in Fig. 3. The full version is notably more expressive, incorporating powerful attention modules [52, 59, 14] and follows a coarse-to-fine matching paradigm [59, 78, 14]. While it delivers more accurate matches it is computationally more expensive (cf. Sec. 5.3). Consequently, we also propose a minimal version of our method that focuses on learning good features for matching, removing the need to learning the matching function itself. Both models comprise a feature extraction module that encodes both the query image and 3D scene points into a feature space, and a matching module that aligns these two feature sets to determine the 2D-3D correspondences. We further detail their architectural designs and supervision in the subsequent subsections.
@@ 5 Experiments
Implementation details.  We use the first two blocks of ConvFormer [74] as the image backbone and initialize it with ImageNet-1K [50] pre-trained weights***The weights can be downloaded from huggingface.co/timm/convformer_b36.sail_in1k_384. We set feature dimensions for coarse and fine matching as $D^{c}=256$ and $D^{f}=128$. For fine matching, we use local window size $w=5$ for image feature cropping. We resize query images to $480\times 480$ for all experiments. We train minimal and full NeRFMatch models using Adam [33] optimizer with canonical learning rate $clr=0.0008/0.0004$ and batch size $cbs=16$ for $30/50$ epochs accordingly. We decay the learning rate based on the cosine annealing scheduling [39]. Our models are trained on $8$ Nvidia V100 GPUs (16/32GB). A MipNeRF model is $5.28$ MB in size. Given a camera pose, it takes $141$ milliseconds to render $6003$D points with its features on a single 16GB Nividia V100 GPU. Our NeRFMatch-Mini and NeRFMatch models are $42.8/50.4$ MB in size and require $37/157ms$ to run a forward pass at $480\times 480$ image resolution. We provide NeRF implementation details in the supplementary material.
5.1 Localization Evaluation
Inference settings.  We use top-1/10 reference poses for Mini/Full NeRFMatch models for outdoor and top-1 for indoor and apply the best pose refinement determined in Sec. 5.4. We provide more details in supplementary.
@@ 5.1 Localization Evaluation
NeRFMatch-Mini NeRF+RGB 19.0/0.3 30.2/0.6 10.3/0.5 11.3/0.4 29.1/0.2 20.0/0.4
NeRFMatch NeRF+RGB 13.0/0.2 19.4/0.4 8.5/0.4 7.9/0.3 17.5/0.1 13.3/0.3
NeRFMatch NeRF 12.7/0.2 20.7/0.4 8.7/0.4 11.3/0.4 19.5/0.1 14.6/0.3
Results on Cambridge Landmarks.  As shown in Tab. 1, our minimal version despite being lightweight, is able to achieve comparative results w.r.t. most of the SCR methods and surpass all APR methods. With a more advanced attention-based hierarchical matching function, our full model achieves the competitive results among all methods. Our experiments fully demonstrate that the NeRF inner features learned via view synthesis are discriminative 3D representations for 2D-3D matching.
Table 2: Indoor localization on 7-Scenes [57]. We report per-scene median rotation and position errors in $(cm,^{\circ})$ and their average across scenes, along with averaged localization recall.
@@
NeRFMatch-Mini NeRF+RGB 1.6/0.5 1.5/0.6 1.4/0.9 3.6/1.0 3.5/0.9 1.7/0.5 8.5/2.1 3.1/0.9 74.4
NeRFMatch NeRF+RGB 0.9/0.3 1.1/0.4 1.4/1.0 3.0/0.8 2.2/0.6 1.0/0.3 9.0/1.5 2.7/0.7 78.2
NeRFMatch NeRF 0.9/0.3 1.1/0.4 1.5/1.0 3.0/0.8 2.2/0.6 1.0/0.3 10.1/1.7 2.8/0.7 78.4
@@
NeRF-only localization.  We further push our method to using NeRF as the only scene representation for localization, which means we would no longer need access to a real image database as part of the scene representation, which requires a significantly bigger storage than a single NeRF model. For this purpose, we propose to perform image retrieval on synthesized images rendered by our NeRF model. This entails a small decrease in performance on Cambridge in translation error. This slight degradation in performance can be attributed to the increased complexity of these scenes for NeRF. This assertion is supported by the observation that, when switching to image retrieval on synthesized images, our model demonstrates almost no change in performance on the indoor 7-Scenes dataset. Our results open the door to a localization pipeline where only a NeRF model is needed.
@@ 5.2 NeRF Feature Ablation
Table 3: NeRF feature ablation on Cambridge [32]. We train NeRFMatch-Mini with different 3D features and compare their localization performance.
@@
Med. Translation ($cm,\downarrow$) 458.0 34.3 28.7 28.4 27.9 28.3 28.3 30.2 61.3
Med. Rotation (${}^{\circ},\downarrow$) 6.5 0.6 0.5 0.5 0.5 0.5 0.5 0.5 1.3
@@ 5.3 NeRFMatch Ablation
- ResNet34 Minimal Per-Scene 32.7/0.6 52.0 32.8 23
- ConvFormer Minimal Per-Scene 34.8/0.6 50.7 42.8 37
NeRFMatch-Mini ConvFormer Minimal Per-Scene 27.9/0.5 59.2 42.8 37
NeRFMatch-Mini (MS) ConvFormer Minimal Multi-Scene 30.8/0.5 53.6 42.8 37
NeRFMatch ConvFormer Full Per-Scene 16.5/0.3 71.3 50.4 157
NeRFMatch (MS) ConvFormer Full Multi-Scene 22.0/0.4 65.2 50.4 157
Training ablation.  In addition to architecture choices, we also examine the influence of different training settings, i.e., training per-scene (default) and training multi-scenes (all scenes within a dataset). Despite NeRF features being trained per-scene, we surprisingly find that both minimal and full NeRFMatch models can be trained to handle multi-scene (MS) localization with only a slight decrease in accuracy w.r.t. the per-scene model. This is a similar finding as recent scene-agnostic SCR methods [72, 63, 62], that extend per-scene SCR to multi-scenes by conditioning SCR on scene-specific 3D points. While scene-agnostic SCR learns to regress directly the 3D features in the form of xyz coordinates, our models learn to find a common ground between image features and NeRF features for matching.
@@ 5.4 Pose Refinement
After conducting matching architecture ablation, we investigate different refinement methods to further increase pose accuracy. We examine two approaches: (i) an iterative approach, where we re-run matching with the last computed estimation as reference pose, and (ii) an optimization approach, where we optimize a reference pose through frozen NeRF weights using a photometric loss and then run the final matching. As shown in Fig. 4, both methods show an improvement over the initial estimate. We assess both the NeRFMatch model and its minimal version (Mini) trained per-scene.
@@
Iterative refinement for computational efficiency.  In the NeRFMatch setting, both refinement approaches show similar early results, since the initial pose estimate is relatively close to the solution, but the iterative approach is more stable over time. Additionally, it is worth noting that the optimization approach incurs a higher computational cost compared to the iterative approach. The runtime for the optimization refinement excluding the matching step (shown in Tab. 4) is $398.4ms$ for a single optimization step and subsequent rendering. In contrast, for a single step of iterative refinement, the runtime amounts to $141.2ms$. Given the better performance and the quicker computational time, we opt for the iterative approach as the default refinement for NeRFMatch.
@@
Optimization refinement for large pose corrections.  In comparison to the iterative method, the optimization approach significantly benefits the minimal matching model. This is because the initial query pose estimation is notably distant compared to the one given by NeRFMatch, and subsequent iterations only result in incremental refinements. The optimization approach takes a more substantial step towards the query pose, and achieving a more optimal reference pose results in improved query pose estimation. We have already seen that the optimization refinement incurs higher computational time, but it also depends on a learning rate schedule, which is sensitive to configuration. Through empirical analysis, we select $1\times 10^{-3}$ as the initial learning rate and apply a cosine annealing learning rate schedule. We set the optimization approach as the default refinement for NeRFMatch-Mini.
Model Best No Refinement Refined Refined
Refinement (top11-1- 1) (top11-1- 1) (top1010-10- 10)
[…]
NeRFMatch-Mini Opt+Match 27.9/0.5/59.2 20.5/0.4/70.9 20.5/0.4/70.9
NeRFMatch Iter. 16.5/0.3/71.3 14.2/0.3/78.2 13.3/0.3/80.8
Figure 4: Refinement ablation on Cambridge Landmarks [32]. On the left side, we depict the average recall for optimization-based (Opt+Match) and iterative (Iter) refinement approaches across multiple iterations. We provide results for both the NeRFMatch and its minimal setting. On the right side, we report averaged median pose error in ($cm/^{\circ}$) and localization recall with the best refinement configurations.